please help me solve 2(4x-2) -4= 64-22x(9-11)
Found 2 solutions by rcmcc, checkley71:
Answer by rcmcc(152)

2(4x-2) -4= 64-22x(9-11)
First step to solving for x is to expand the brackets
8x-4-4=64-44x
simplify the equation
8x-8=64-44X
move the x's to one side.
52x-8=64
now move the numbers to one side
52x=72
now divide by 52 from each side to isolate x

x=(18/13)

Answer by checkley71(8403)

2(4X-2)-4=64-22X(9-11)
8X-4-4=64+22X*-2
8X-44X=64+4+4
-36X=72
X=72/-36
X=-2
PROOF
2(4*-2-2)-4=64-22*-2(9-11)
2(-8-2)-4=64+44(-2)
2(-10)-4=64-88
-20-4=-24
-24=-24